OpenStudy (anonymous):

Can someone please solve and explain this word problem? "A rectangular painting measures 12 inches by 16 inches and is surrounded by a frame of uniform width around the four edges. The perimeter of the rectangle formed by the painting and its frame is 72 inches. Determine the width of the frame."

OpenStudy (amistre64):

this is the pool problem again

OpenStudy (amistre64):

x+12+x = width x+16+x = length

OpenStudy (amistre64):

perimeter is 72 72/2 = 36 hmmm

OpenStudy (amistre64):

nah, just add up twice the length and width and equal it out to 72

OpenStudy (amistre64):

4x+24+4x+32 = 72 if i see it right; where x = width of frame

OpenStudy (mertsj):